# EarthMind: Towards Multi-Granular and Multi-Sensor Earth Observation with Large Multimodal Models

### v1

@@ 5.4 Ablations
EarthMind significantly enhances multi-granularity and multi-sensor EO understanding through the proposed (i) Spatial Attention Prompting (SAP) and (ii) Cross-modal Fusion. We delve into each of these components in the following sections. More discussions can be seen in Appendix G.
Effectiveness of SAP. First, we evaluate the effectiveness of our proposed SAP method. To ensure a fair comparison, we train models from scratch on the training splits of RefSegRS and RRSIS-D, and report results on their corresponding test sets. As shown in Tab. 5, incorporating SAP significantly improves segmentation performance. Second, we visualize the attention maps between the <SEG> token and image tokens across transformer layers within the LLM. As illustrated in Fig. 4, SAP effectively suppresses attention drift and helps reallocate attention toward the anchor regions corresponding to the query. Lastly, we investigate how the performance of SAP varies with the supervision layer. Results show that applying supervision at middle layers yields the strongest performance boost, as these layers balance low-level spatial detail and high-level semantics. Importantly, SAP introduces no additional parameters, making it a lightweight yet effective enhancement for pixel-level grounding within LLMs.
@@
Effectiveness of Cross-modal Fusion. To evaluate the effectiveness of our cross-modal fusion design, we conduct ablations on the multi-sensor setting with three configurations: (1) w/o Modality Alignment; (2) Naive Concatenate: visual tokens from different modalities are directly concatenated and passed to the LLM; (3) Naive Attention: token importance is computed using cosine similarity between paired SAR and optical features. We report both multiple-choice accuracy and referring expression segmentation results on EarthMind-Bench. As shown in Tab. 5, our full model significantly outperforms all baselines, demonstrating the importance of both alignment and adaptive weighting for robust fusion. Furthermore, to assess the informativeness of modality-selected tokens, we compare different token retention strategies. For baselines, we randomly discard a portion of the input tokens before feeding them into the LLM. In contrast, our method selectively drops tokens with lower learned importance weights. As shown in Fig. 4, even under the same retention ratio, EarthMind retains more informative content, leading to better performance. This demonstrates that our method reserves the most complementary information across them.
6 Limitations and Conclusion
Limitations of EarthMind. Training EarthMind demands considerable computational resources due to its use of multiple visual encoders for multi-level understanding. A promising direction is to optimize the architecture via Mixture-of-Experts or knowledge distillation to reduce redundancy. Additionally, a modality-aligned encoder that jointly embeds heterogeneous sensor inputs into a shared semantic space could further improve efficiency.
In this work, we propose EarthMind, a unified framework for multi-granular and multi-sensor understanding in Earth Observation (EO). To enhance pixel-level performance, we introduce Spatial Attention Prompting (SAP), which guides attention reallocation within the LLM. We also design a Cross-modal Fusion mechanism that aligns heterogeneous modalities into a shared space and adaptively selects informative features. To support evaluation, we curate EarthMind-Bench, the first benchmark tailored for multi-sensor fusion in LMMs. Extensive experiments validate the effectiveness and strong generalization of EarthMind across diverse EO tasks.
@@ Appendix A Overview of Appendix
  • B: Discussions and Borader Impact.
  • C: Details of EarthMind.
  • D: Details of EarthMind-Bench.
  • E: Experimental Settings.
  • F: Training Dataset.
  • G: More Ablation Studies.
  • H: More Visualization Results.
Appendix B Discussions and Broader Impact
We discuss the limitation of EarthMind in Sec. 6, here we further discuss the limitation of EarthMind-Bench. Currently, EarthMind-Bench includes only paired optical (RGB) and SAR modalities. To enable a more comprehensive evaluation of LMMs’ multi-sensor fusion capabilities, future extensions should incorporate additional sensing modalities such as multispectral, hyperspectral, and infrared imagery. Moreover, benchmarking diverse fusion scenarios involving more than two modalities would further reflect EO challenges.
@@ Appendix B Discussions and Broader Impact
Broader Impact. EarthMind offers a broader impact to the large multimodal model (LMM) community in both methodology and application. On the one hand, it demonstrates how vision-language models can be extended to handle multi-granular tasks via the proposed Spatial Attention Prompting mechanism, which reallocates LLM attention to task-relevant regions. This architectural design is not limited to Earth Observation (EO), but can be generalized to other domains requiring fine-grained spatial reasoning, such as medical imaging and autonomous driving.
On the other hand, for the EO community, EarthMind contributes both algorithmically and empirically: it introduces an effective multi-sensor fusion framework and proposes a scalable training pipeline, accompanied by a new benchmark and instruction-tuning dataset. These efforts can benefit various downstream applications in remote sensing, including disaster forecasting, infrastructure monitoring, and route navigation.
Appendix C Details of EarthMind
@@
EarthMind is built upon the InternVL-2 framework [49]. In InternVL-2, each image is divided into multiple patches at pre-defined scales. Each patch is processed by the visual encoder and encoded into 256 tokens. For instance, an image with 4 patches (plus a global image token) yields $(4+1)\times 256$ visual tokens in total. For region-level perception, we adopt the region encoder from GPT4RoI [66], where each detected region is encoded into 256 tokens as well. To support multi-sensor input, non-optical imagery (e.g., SAR or multispectral data) is transformed into a synthetic video-like sequence by stacking pre-processed frames. This sequence is then concatenated before the input query, following the protocol of multi-frame vision-language models.
@@
Figure 5: Detailed data statistics of the EarthMind-Bench.
Appendix D Details of EarthMind-Bench
Overview. We present a more detailed analysis of EarthMind-Bench in Fig. 5. Subfigures (a–c) illustrate the distribution of task categories, while (d) and (e) report the number of samples per task. Notably, our benchmark includes 438 referring expression segmentation samples paired with corresponding binary masks, highlighting its support for fine-grained pixel-level grounding. We further visualize the word clouds of questions (f) and answers (g), showing that EarthMind-Bench covers a wide variety of object types and semantics, enabling comprehensive evaluation across perception and reasoning tasks.
@@ Appendix D Details of EarthMind-Bench
  • 1. Object Existence Determine whether a specific object or class exists in the image (e.g., Is there a bridge in the image?”). Guideline: The object should be clearly visible and occupy a non-trivial area. If heavily occluded or ambiguous, mark as Not present.”
@@
  • 2. Hallucination Detection Detect whether the model falsely recognizes objects that do not exist. Guideline: Ask for fine-grained distinctions between semantically similar objects (e.g., “Is there a train or a bus in the image?”). Confirm that the mistaken object does not exist even under partial occlusion.
@@
  • 3. Object Counting Count the number of objects from a given category (e.g., “How many buildings are visible in the image?”). Guideline: Only count objects that are visually distinguishable and not clustered into ambiguous shapes. Accept a small margin of error (±1) in complex scenes.
@@
  • 4. Spatial Relationship Identify relative positions between objects (e.g., What is next to the waterbody?”). Guideline: Use cardinal or contextual spatial relations (e.g., to the left of,” adjacent to,” surrounded by”). Annotators should verify object co-existence and relative proximity.
@@
  • 5. Route Planning Generate or select feasible navigation routes from a start to a target location. Guideline: Ensure the proposed path avoids obstacles (e.g., rivers, buildings), follows terrain constraints (e.g., avoid steep hills), and adheres to logical movement (e.g., roads preferred over fields).
@@
  • 6. Image Captioning Generate descriptive sentences summarizing the content and layout of the image. Guideline: Captions should mention key objects, land types, spatial layout, and human-made structures if visible. Avoid hallucination and be concise yet informative.
@@
  • 7. Disaster Forecasting Assess the likelihood of disaster based on visual evidence (e.g., “Is this area prone to flooding?”). Guideline: Use cues such as terrain (lowlands), proximity to water, lack of infrastructure, or signs of previous disaster impact. Avoid speculative answers.
@@
  • 8. Urban Development Assessment Evaluate whether an area is suitable for urban development. Guideline: Consider factors such as flat terrain, absence of natural barriers, existing road access, and land cover type. Annotators should justify suitability with visual cues.
@@
Evaluation Metrics. For multiple-choice tasks, we evaluate model performance using standard accuracy, measuring the percentage of predictions that exactly match the ground-truth option. For open-ended tasks, inspired by [67], we adopt a GPT-4-based evaluation protocol that assesses the alignment between the model-generated answers and human annotations. As illustrated in Fig. 6, we prompt GPT-4 to score the responses based on semantic similarity and correctness, following a structured rubric to ensure consistent evaluation across tasks.
Figure 6: The illustration of open-ended task evaluation prompt.
Appendix E Details of Experimental Settings
@@
We elaborate on the training and inference details of EarthMind. Specifically, we report the hyperparameters in the fine-tuning stage, as shown in Tab. 6.
@@
Table 6: Hyperparameters of EarthMind.
Appendix F Training Dataset
EarthMind is trained on large-scale natural image datasets, including LLaVA-665K [6], 56K referring expression data [68, 69], and 214K grounding conversation generation samples [9]. These datasets cover image-level captioning, VQA, and text-driven segmentation. We also incorporate 724K region-level descriptions from the Osprey dataset [70] to improve region-level understanding capacity. Second, we introduce EO-specific multimodal data to adapt EarthMind to the remote sensing domain. This includes 1M VQA data from EarthGPT [17], 142K EO conversations from VRSBench [35], 31K region-level captions from DIOR-RSVG [54], and 21K referring segmentation samples from RRSIS-D [56] and RefSegRS [55]. Moreover, we involve 500k multi-spectral data, including BigEarthNet [57] and SoSAT-LCZ42 [57]. Third, we utilize our synthesized multi-sensor conversation corpus (20K RGB-SAR paired dialogues) and selectively retain examples from earlier stages to mitigate catastrophic forgetting.
Paired Multi-Sensor Training Data Curation. We construct a high-quality paired RGB–SAR training corpus from five publicly available datasets: OpenEarthMap-SAR [44], DFC2023 Track2 [45], WHU-OPT-SAR [46], MSAW [47], and MultiResSAR [48]. All RGB–SAR pairs are sampled from their training splits to avoid overlap with EarthMind-Bench. To scale data generation, we design an automatic synthetic pipeline based on GPT-4o. Since GPT-4o performs better on RGB imagery than SAR, we use RGB images and their classification labels as input prompts. The prompt is tailored for large vision-language models (VLMs), comprising the image and a list of detected object categories (e.g., buildings, roads, water bodies). The model is first asked to generate a comprehensive caption summarizing the scene and spatial relationships among objects. Then, it is instructed to synthesize 3–10 diverse question–answer pairs. These cover various reasoning types, such as: (1) Object Existence (e.g., “Is there a river in the image?”), (2) Counting (e.g., “How many buildings are visible?”), (3) Spatial Relation (e.g., “What is next to the forest?”), (4) Object Localization (e.g., “Where is the road located?”), and (5) Scene-Level Understanding (e.g., “Is this area suitable for agriculture?”). All outputs are returned in a structured JSON format for seamless integration into our training pipeline. In total, we curate 20k synthetic multi-sensor samples, which serve as paired RGB–SAR instruction data for adaptation of multi-sensor LMMs.
Table 7: Ablations on the non-rgb data processing.

### v2

@@ 5.4 Ablations
EarthMind significantly enhances multi-sensor EO data understanding through the proposed Hierarchical Cross-modal Attention (HCA) module and the curated FusionEO dataset. We examine each component below, with additional analysis in Appendix H.
@@
The effectiveness of HCA. To evaluate the effectiveness of our HCA module, we conduct ablations on the multi-sensor setting by comparing with the better single-modality performance (optical or SAR) and three fusion configurations: (1) Concatenate: visual tokens from different modalities are directly concatenated and passed to the LLM; (2) Element-wise Averaging: visual tokens are summed and averaged to obtain fused representations; (3) Naive Attention: token importance is computed using cosine similarity between paired SAR and optical features. We report both multiple-choice task performance and referring expression segmentation results on EarthMind-Bench.
As shown in Tab. 4, our method significantly outperforms all naive fusion methods, especially on challenging tasks like object counting and spatial relationship reasoning. These fine-grained tasks require detailed visual perception and reasoning, making them more susceptible to poor optical image quality and thus benefiting from complementary information provided by SAR imagery. To understand the underlying mechanism, we calculate the Modality Attention Score (MAS) proposed in Eq. 1 and visualize the results for HCA and naive concatenation across different layers. The analysis reveals a dramatic attention imbalance in naive concatenation: optical tokens receive significantly higher attention than SAR tokens, creating a substantial attention gap throughout all layers. This confirms that the model is nearly blind to SAR information. In contrast, our HCA method significantly reduces this attention gap, achieving a more balanced distribution. This rebalancing is particularly pronounced in shallow layers where fundamental perceptual representations are established, enabling the model to effectively leverage complementary information from both modalities.
We also visualize a typical case in Fig. 4 (a), which reveals interesting findings. First, low-quality regions in optical images receive low attention weights, where SAR can provide complementary information. Moreover, the attention weights are highly relevant to the query: for global tasks like classification, the model primarily relies on the optical modality for scene understanding, while for fine-grained tasks like segmentation, SAR modality gains increased importance, especially in query-relevant regions. For instance, when segmenting roads, the model learns to leverage complementary information from both modalities, with SAR providing crucial structural details that may be obscured or ambiguous in optical imagery. We further calculate the task-specific fusion weights using Eq. 7 and visualize the results in Fig. 4 (b) to elaborate this phenomenon, where optical modality dominates in semantic-rich tasks while SAR modality begins to contribute significantly in fine-grained tasks.
Figure 4: (a) Visualization of learned attention weights of optical and SAR under the different queries. (b) Statistics of attention weights in different tasks. (c) The ablations of the proposed FusionEO data.
The effectiveness of FusionEO. Our proposed FusionEO employs a three-stage curation pipeline. To validate its effectiveness, we report the performance scaling on EarthMind-Bench in Fig. 4 (c). As shown, our curated data achieves consistent performance gains with increasing data scale. Compared to Stage 1 data which contains coarse-grained information, the RoI information injection in subsequent stages provides more fine-grained supervision. Building upon this foundation, our final dataset generates diversified QA pairs, thus achieving the best generalization performance.
6 Conclusion
In this work, we propose EarthMind, a unified framework for multi-sensor Earth Observation data understanding, specifically designed for optical and SAR data fusion. We introduce the Hierarchical Cross-modal Attention (HCA) module, which adaptively selects task-related cross-sensor representations for enhanced LLM reasoning. EarthMind handles diverse EO tasks from image-level understanding to pixel-level segmentation. We develop FusionEO, a 20K paired instruction-tuning dataset, and curate EarthMind-Bench, the first benchmark for multi-sensor fusion in MLLMs. Extensive experiments demonstrate that EarthMind achieves state-of-the-art performance while consistently outperforming existing MLLMs across diverse EO benchmarks, validating its effectiveness and strong generalization.
@@ Appendix A Overview of Appendix
  • B: The Use of Large Language Models (LLMs).
  • C: Discussions and Broader Impact.
  • D: Details of EarthMind.
  • E: Details of EarthMind-Bench.
  • F: Experimental Settings.
  • G: Training Dataset.
  • H: More Ablation Studies.
  • I: More Visualization Results.
Appendix B The Use of Large Language Models (LLMs)
In accordance with ICLR policies on LLM usage, we disclose that no Large Language Models were used in any aspect of this research work. All content, including research conception, methodology development, experimental design, code implementation, data analysis, result interpretation, and manuscript preparation, was conducted entirely by the authors without LLM assistance. The authors assume full responsibility for all content, claims, and conclusions in this paper.
Appendix C Discussions and Broader Impact
In this section, we illustrate the limitations, the distinction between our method and existing ones, as well as the broader impact.
Limitation and Future Work. Training EarthMind requires considerable computational resources due to its use of multiple visual encoders for multi-level understanding. A promising direction is to optimize the architecture via Mixture-of-Experts or knowledge distillation to reduce redundancy. Additionally, a modality-aligned encoder that jointly embeds heterogeneous sensor inputs into a shared semantic space could further improve efficiency. Moreover, our EarthMind-Bench currently contains only optical-SAR data; future work will incorporate additional sensing modalities such as multispectral, hyperspectral, and infrared imagery. Furthermore, benchmarking diverse fusion scenarios involving more than two modalities would better reflect real-world EO challenges.
Comparison with Existing Works. The key distinction of our work from existing models is that EarthMind is the first to address both multi-sensor fusion and multi-granular tasks simultaneously. For multi-sensor capability, our model can flexibly support single or multiple visual inputs; for multi-granular understanding, it can handle multiple levels of granularity, from pixel-level segmentation and region-level semantic understanding to image-level scene classification. As summarized in Tab. A, achieving fine-grained, multi-sensor comprehension in EO remains largely unresolved by existing approaches.
@@ Appendix C Discussions and Broader Impact
Broader Impact. EarthMind offers significant impact to both the large multimodal model (LMM) community and practical applications. Methodologically, it demonstrates how vision-language models can be extended to handle multi-granular tasks via the proposed Hierarchical Cross-modal Attention mechanism, which adaptively allocates attention to task-relevant regions and modalities. This architectural design extends beyond Earth Observation and can be generalized to other domains requiring fine-grained spatial reasoning, such as medical imaging and autonomous driving. For the EO community, EarthMind contributes both algorithmically and empirically: it introduces an effective multi-sensor fusion framework and proposes a scalable training pipeline, accompanied by a comprehensive benchmark and instruction-tuning dataset. These contributions can benefit various downstream applications in remote sensing, including disaster forecasting, infrastructure monitoring, and environmental assessment.
Table A: (Left) Comparison of EarthMind with existing EO LMMs. EarthMind supports both multi-granular and multi-sensor understanding. (Right) Comparison of EarthMind-Bench with existing EO benchmarks in terms of multi-sensor support, granularity, and task level. “S” denotes single-modality input; “M” indicates multiple modalities (used independently); “F” represents paired sensor fusion. “MC” and “OE” refer to multiple-choice and open-ended formats, respectively.
Method Multi-Granular Multi-Sensor Image Region Pixel Handling Fusion RSGPT (Hu et al., 2025) GeoChat (Kuckreja et al., 2024) EarthGPT (Zhang et al., 2024b) Earthmarker (Zhang et al., 2024a) LHRS-bot (Muhtar et al., 2024) SkyEyeGPT (Zhan et al., 2025) Skysensegpt (Zhan et al., 2025) EarthDial (Soni et al., 2024) GeoPixel (Shabbir et al., 2025) RSUniVLM (Liu & Lian, 2024) EarthMind
Benchmark Multi Sensor Multi Gran. Multi Level Task Type RSIEval (Hu et al., 2025) S MC + OE HnstD (Pang et al., 2025) S MC + OE GeoChat-Bench (Kuckreja et al., 2024) S OE VRSBench (Li et al., 2024b) S MC + OE LHRS-Bench (Muhtar et al., 2024) S MC VLEO-Bench (Zhang & Wang, 2024) S MC + OE FIT-RSRC (Luo et al., 2024) S MC UrBench (Zhou et al., 2025) S MC XLRS-Bench (Wang et al., 2025) S MC + OE GEOBench-VLM (Danish et al., 2024) M MC + OE EarthMind-Bench M + F MC + OE
Appendix D Details of EarthMind
@@
EarthMind is built upon the InternVL-2 framework (Chen et al., 2024b). In InternVL-2, each image is divided into multiple patches at pre-defined scales. Each patch is processed by the visual encoder and encoded into 256 tokens. For instance, an image with 4 patches (plus a global image token) yields $(4+1)\times 256$ visual tokens in total. To support multi-sensor input, non-optical imagery (e.g., SAR or multispectral data) is transformed into a synthetic video-like sequence by stacking pre-processed frames. This sequence is then concatenated before the input query, following the protocol of multi-frame vision-language models.
@@
Figure A: Detailed data statistics of the EarthMind-Bench.
Appendix E Details of EarthMind-Bench
Overview. We present a more detailed analysis of EarthMind-Bench in Fig. A. Subfigures (a–c) illustrate the distribution of task categories, while (d) and (e) report the number of samples per task. Notably, our benchmark includes 438 referring expression segmentation samples paired with corresponding binary masks, highlighting its support for fine-grained pixel-level grounding. We further visualize the word clouds of questions (f) and answers (g), showing that EarthMind-Bench covers a wide variety of object types and semantics, enabling comprehensive evaluation across perception and reasoning tasks.
@@ Appendix E Details of EarthMind-Bench
  • 1. Object Existence. Determine whether a specific object or class exists in the image (e.g., Is there a bridge in the image?”). Guideline: The object should be clearly visible and occupy a non-trivial area. If heavily occluded or ambiguous, mark as Not present.”
@@
  • 2. Hallucination Detection. Detect whether the model falsely recognizes objects that do not exist. Guideline: Ask for fine-grained distinctions between semantically similar objects (e.g., “Is there a train or a bus in the image?”). Confirm that the mistaken object does not exist even under partial occlusion.
@@
  • 3. Object Counting. Count the number of objects from a given category (e.g., “How many buildings are visible in the image?”). Guideline: Only count objects that are visually distinguishable and not clustered into ambiguous shapes. Accept a small margin of error (±1) in complex scenes.
@@
  • 4. Spatial Relationship. Identify relative positions between objects (e.g., What is next to the waterbody?”). Guideline: Use cardinal or contextual spatial relations (e.g., to the left of,” adjacent to,” surrounded by”). Annotators should verify object co-existence and relative proximity.
@@
  • 5. Route Planning. Generate or select feasible navigation routes from a start to a target location. Guideline: Ensure the proposed path avoids obstacles (e.g., rivers, buildings), follows terrain constraints (e.g., avoid steep hills), and adheres to logical movement (e.g., roads preferred over fields).
@@
  • 6. Image Captioning. Generate descriptive sentences summarizing the content and layout of the image. Guideline: Captions should mention key objects, land types, spatial layout, and human-made structures if visible. Avoid hallucination and be concise yet informative.
@@
  • 7. Disaster Forecasting. Assess the likelihood of disaster based on visual evidence (e.g., “Is this area prone to flooding?”). Guideline: Use cues such as terrain (lowlands), proximity to water, lack of infrastructure, or signs of previous disaster impact. Avoid speculative answers.
@@
  • 8. Urban Development Assessment. Evaluate whether an area is suitable for urban development. Guideline: Consider factors such as flat terrain, absence of natural barriers, existing road access, and land cover type. Annotators should justify suitability with visual cues.
@@
Evaluation Metrics. For multiple-choice tasks, we evaluate model performance using standard accuracy, measuring the percentage of predictions that exactly match the ground-truth option. For open-ended tasks, inspired by (Zhou et al., 2024), we adopt a GPT-4-based evaluation protocol that assesses the alignment between the model-generated answers and human annotations. As illustrated in Fig. B, we prompt GPT-4 to score the responses based on semantic similarity and correctness, following a structured rubric to ensure consistent evaluation across tasks.
Figure B: The illustration of open-ended task evaluation prompt.
Appendix F Details of Experimental Settings
@@
We elaborate on the training and inference details of EarthMind. Specifically, we report the hyperparameters in the fine-tuning stage, as shown in Tab. B.
@@
Table B: Hyperparameters of EarthMind.
Figure C: The pipeline of FusionEO curation.
Appendix G Training Dataset
EarthMind is trained on large-scale natural image datasets, including LLaVA-665K (Liu et al., 2023a), 56K referring expression data (Kazemzadeh et al., 2014; Yu et al., 2016), and 214K grounding conversation generation samples (Rasheed et al., 2024). These datasets cover image-level captioning, VQA, and text-driven segmentation. We also incorporate 724K region-level descriptions from the Osprey dataset (Yuan et al., 2024a) to improve region-level understanding capacity. Second, we introduce EO-specific multimodal data to adapt EarthMind to the remote sensing domain. This includes 1M VQA data from EarthGPT (Zhang et al., 2024b), 142K EO conversations from VRSBench (Li et al., 2024b), 31K region-level captions from DIOR-RSVG (Zhan et al., 2023), and 21K referring segmentation samples from RRSIS-D (Liu et al., 2024b) and RefSegRS (Yuan et al., 2024b). Moreover, we involve 500k multi-spectral data, including BigEarthNet (Sumbul et al., 2019) and SoSAT-LCZ42 (Sumbul et al., 2019). Third, we utilize our synthesized multi-sensor conversation corpus (20K RGB-SAR paired dialogues) and selectively retain examples from earlier stages to mitigate catastrophic forgetting.
Paired Multi-Sensor Training Data Curation. We construct a high-quality paired optical-SAR training corpus from six publicly available datasets: BigEarthNet-MM (Sumbul et al., 2021), OpenEarthMap-SAR (Xia et al., 2025), DFC2023 Track2 (Persello et al., 2023), WHU-OPT-SAR (Li et al., 2022), MSAW (Shermeyer et al., 2020), and MultiResSAR (Zhang et al., 2025b). All optical-SAR pairs are sampled from their training splits to avoid overlap with EarthMind-Bench.
To scale data generation, we design an automatic three-stage synthetic pipeline leveraging GPT-4o, as shown in Fig. C: Stage 1: Metadata Preparation involves extracting modality information (optical and SAR) and category labels (e.g., rangeland, road, building) for each image pair. Stage 2: RoI-based Summarization uses the optical image and metadata as input to GPT-4o to generate comprehensive scene descriptions. The model produces both short captions summarizing key objects and spatial relationships, and detailed long captions that provide fine-grained descriptions of regions of interest and their contextual relationships. Stage 3: Self-instruct VQA Generation creates diverse question-answer pairs based on the generated captions and multi-sensor context. These cover various reasoning types including: (1) Object Existence (e.g., ”Is there a river in the image?”), (2) Counting (e.g., ”How many buildings are visible?”), (3) Spatial Relations (e.g., ”What is next to the forest?”), (4) Object Localization (e.g., ”Where is the road located?”), and (5) Scene-Level Understanding (e.g., ”Is this area suitable for agriculture?”).
All outputs are structured in JSON format for seamless integration. This three-stage approach ensures that the generated instruction data captures both coarse-grained scene understanding and fine-grained spatial reasoning across modalities. In total, we curate 30K synthetic multi-sensor samples as our FusionEO dataset for training cross-sensor MLLMs.
Table C: Ablations on the non-rgb data processing.